. HI decomposes according to the equation 2 HI (g) → H2 (g) + I2 (g). The rate constant k for this reaction was measured at several different absolute temperatures T, and an Arrhenius plot of ln k vs. 1/T had a slope of -2.24 × 10^4 K. What is the activation energy Ea for this reaction?

A. 1.86 × 105 J/mol

B. 3.28 × 103 J/mol

C. -1.86 × 105 J/mol

D. -3.28 × 103 J/mol
Answer #1

use,
lnk = lnA - Ea/(R*T)
so,
when we plot graph between ln k and 1/T
slope will be (-Ea/R)
so,
-Ea/R = -2.24*10^4
Ea/8.314 = 2.24*10^4
Ea = (1.86*10^5) J/mol

Answer : option A